# The interplay of WGC and WCCC via charged scalar field fluxes in the RPST framework

### v1

@@ 7 Conclusions
We employed the restricted phase space(RPS) thermodynamics and investigated the weak cosmic censorship conjecture (WCCC). Also in this paper, we used the equivalence of mass and energy principle and discussed the relationship between two important conjectures as WCCC which is derived by general relativity and WGC is derived from string theory. Here, we assumed that the internal charge and energy of the black hole change with the superradiance and absorption of the particle (or scalar field) in an infinitesimal time interval. By using the reproduced first law of thermodynamics, we have shown that the second law holds in conditions where the temperature has a lower limit as $(T\geq\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)})$. But the second law of thermodynamics can be violated in conditions where the temperature has an upper limit and the black hole is in equilibrium $\omega=\hat{q}\hat{\Phi_{h}}$. Also, here we point out that when the central charge does not change, the obtained results are the same as the results obtained in normal space. In addition, we showed using the RPS thermodynamics that the WCCC holds despite the input and output flux in the extremal state and close to the extremal state. Therefore, the result obtained in the RPS is consistent with the results obtained in the normal space and the extended phase space(EPS). By applying the equivalence mass and energy principle for the particle and second-order approximation in the state close to the extreme black hole, we find that when we have superradiance , particles are emitted that obey the WGC, of course, with the condition that the $C>>\hat{Q}$. In this case, the black hole will be further away from its extreme state. Also, a black hole will only attract a particle that obeys the WGC if the black hole is very small and the central charge is larger than the scaled electric charge $C>>\hat{Q}$. In this case, the black hole becomes closer to its extreme state.
8 Acknowledgments
The authors would like to express their gratitude to Saeed Noori Gashti for his insightful discussions.

### v2

@@ 7 Conclusions
We employed the RPS thermodynamics and investigated the WCCC. Also in this paper, we used the equivalence of mass and energy principle and discussed the relationship between two important conjectures WCCC which is derived from general relativity, and WGC is derived from string theory. Here, we assumed that the internal charge and energy of the black hole change with the superradiance and absorption of the particle (or scalar field) in an infinitesimal time interval. By using the reproduced first law of thermodynamics, we have shown that the second law holds in conditions where the temperature has a lower limit as $(T\geq\frac{1}{2\pi r_{h}}\bigg{(}1+\frac{r_{h}^{2}}{\ell^{2}}\bigg{)})$. But the second law of thermodynamics can be violated in conditions where the temperature has an upper limit and the black hole is in equilibrium $\omega=\hat{q}\hat{\Phi_{h}}$. Also, here we point out that when the central charge does not change, the obtained results are the same as the results obtained in normal space. In addition, we showed using the RPS thermodynamics that the WCCC holds despite the input and output flux in the extremal state and close to the extremal state. Therefore, the result obtained in the RPS is consistent with the results obtained in the normal space and the extended phase space(EPS). By applying the equivalence mass and energy principle for the particle and second-order approximation in the state close to the extreme black hole, we find that when we have superradiance, particles are emitted that obey the WGC, of course, with the condition that the $C>>\hat{Q}$. In this case, the black hole will be further away from its extreme state. Also, a black hole will only attract a particle that obeys the WGC if the black hole is very small and the central charge is larger than the scaled electric charge $C>>\hat{Q}$. In this case, the black hole becomes closer to its extreme state.